40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11917     Entry time: Thu Jan 7 04:28:39 2016     In reply to: 11912     Reply to this: 11919
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX X Freq Noise measured 

[ericq, Gautam]

Brief summary of tonights work:

  • Locked Marconi to AUX X vs PSL beat at around 320MHz, PSL shutter closed (i.e. both lasers free running)
  • Measured control signal spectrum at various laser diode currents, crystal temperatures. Oddly, spectra remained consistent across these variables. 
  • Measured OLG of PLL to calibrate into open-loop frequency noise of the beat, found UGF ~30kHz

Our "requirement" for the end laser is as follows: We expect to (and have in the past) achieved ALS sensitivity of 1Hz/rtHz at 100 Hz. If the end PDH loop is 1/f from 100Hz-10kHz, then we have 40dB of supression at 100Hz, meaning the free running AUX laser noise should be no more than 100Hz/rtHz at 100Hz.

So, if we expect both the PSL and AUX lasers to have this performance when free running, we would get the green curve below. We do not. frown


I'll post more details about the exact currents, temperatures and include calibrated plots for the >30kHz range later. Here's the OLG for kicks. 

Attachment 1: PLLspec.pdf  24 kB  | Hide | Hide all | Show all
PLLspec.pdf
Attachment 2: PLL_OLG.pdf  22 kB  | Show | Hide all | Show all
ELOG V3.1.3-